OpenStudy (lgbasallote):

i always wanted to know how to solve these: a particle moves according to the equation \(x = 10t^2\), where x is in meters and t is in seconds (a) Find the average velocity for the time interval from 2.00 s to 3.00 s

ganeshie8 (ganeshie8):

tale integral of x from 2 to 3

OpenStudy (lgbasallote):

what about with no calclulus involved?

OpenStudy (shubhamsrg):

find x for t=2 and 3 calculate the difference divide by time interval (here, 1 sec) ans is diff/time

OpenStudy (lgbasallote):

uhh would you do it step by step not flood the steps into me coz i cant visualize that o.O

OpenStudy (lgbasallote):

so x = 40 and x = 90

OpenStudy (lgbasallote):

dofference is 50

OpenStudy (lgbasallote):

so the answer is 50?

OpenStudy (shubhamsrg):

yep..so far so good.

OpenStudy (shubhamsrg):

seems right to me

OpenStudy (anonymous):

it's all you, shubham.

OpenStudy (lgbasallote):

can you tell me the reasoning for this solution?

OpenStudy (shubhamsrg):

x denotes distance..or infact position.. so you get to know that at t=2,,position was as 40 and at 3,,it was at 90 so it travelled 50 metres in between.. which gived you the ans

ganeshie8 (ganeshie8):

humm average velocity should be 10 m/s

OpenStudy (lgbasallote):

x denotes distance....then does that mean 10 is speed?

OpenStudy (anonymous):

no, x purely informs us about the position of the object in relation to the chosen origin.

OpenStudy (shubhamsrg):

why 10 ? o.O

ganeshie8 (ganeshie8):

x = 10t^2 v = 20t between 2 and 3 average speed = (20 x 3 - 20 x2) / 2 = 10 m/s calculus gives clean answers for these...

OpenStudy (lgbasallote):

so how come the one in between the intervals gives the average velocity?

mathslover (mathslover):

a particle moves according to the equation x=10t2, where x is in meters and t is in seconds (a) Find the average velocity for the time interval from 2.00 s to 3.00 s since x is in meters that means it denotes distance/displacement .. and t denotes time ............. this is the case of average instantaneous velocity

OpenStudy (anonymous):

We need Feynman to establish a consensus here.

OpenStudy (lgbasallote):

let's not introduce some foreigners @Limitless im xenophibic ;)

OpenStudy (shubhamsrg):

well v you obtain by differentiating,,is instantaneous and finding mean of the 2 inst vel. i guess wont be right,,i may be wrong,,please correct me if so..

OpenStudy (lgbasallote):

why is it instantaneous velocity again @mathslover ?

OpenStudy (anonymous):

so, if you take the difference between the positions of the object at two different times, we get the displacement of that object.

OpenStudy (lgbasallote):

^right..so how did it become velocity

OpenStudy (lgbasallote):

90 and 40 are positions right?

OpenStudy (anonymous):

so wat is the qn............x is distance travelled in that instant r x is the position at time t :/

OpenStudy (anonymous):

avg displacement/time=avg velocity

OpenStudy (anonymous):

\[x(t)=10t^2\] \[v_{\text{avg}}=\frac{v_{\text{i}}+v_{\text{f}}}{2}\]

OpenStudy (lgbasallote):

ohh the time is 3 - 2 @rajathsbhat ?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

That was also what I was concluding, Rajath.

OpenStudy (anonymous):

Because the interval is \(2\) to \(3\), which leaves room for only \(1\) second.

OpenStudy (lgbasallote):

i see..makes sense..well a follow-up question is to find the average velocity fo tehe time interval from 2.00 to 2.10 s will it be \[\frac{10(2.10)^2 - 10(2)^2}{2.10-2}?\]

mathslover (mathslover):

Find the average velocity for the time interval from 2.0 s to 3.0 s. x(2 s) = 10 (2)2 = 10 (4) = 40 m x(3 s) = 10 (3)2 = 10 (9) = 90 m \[x = x_f - x_i = x(3 s) - x(2 s) = 90 m - 40 m = 50 m\] t = 1.0 s \[v_avg = x / t = 50 m / 1.0 s = 50 m/s\]

OpenStudy (anonymous):

@Limitless' equation for avg velocity is correct as long a the acceleration is constant.

OpenStudy (anonymous):

Correct, @rajathsbhat. Aren't we operating under that assumption? I apologize if that is not the case.

mathslover (mathslover):

the last equation is \[v_{avg.}=\frac{x}{t}=\frac{50ms^{-1}}{1.0s}=50ms^{-1}\]

OpenStudy (anonymous):

yes, we are.

OpenStudy (lgbasallote):

^wait what?!

OpenStudy (anonymous):

That's incorrect, @mathslover . That is the velocity, not the average velocity.

OpenStudy (lgbasallote):

where did -1 come from???

OpenStudy (lgbasallote):

and will someone look at my solution for the follow-up question lol

OpenStudy (anonymous):

plus, the unit of x is not ms-1

OpenStudy (lgbasallote):

what is this ms-1?! o.O

mathslover (mathslover):

ms-1 is meter per second...

OpenStudy (anonymous):

\[ms ^{-1}\]

OpenStudy (lgbasallote):

oh..a fancy way of writing m/s lo

OpenStudy (lgbasallote):

lol*

OpenStudy (anonymous):

lgba, what have you used in ur follow-up sol?

OpenStudy (lgbasallote):

2.10 and 2?

OpenStudy (anonymous):

no, the numerator.

OpenStudy (anonymous):

Wouldn't \[v_{avg}=\frac{\frac{x(t_1)}{t_1}+\frac{x(t_2)}{t_2}}{2}\] be a straightforward formula?

OpenStudy (lgbasallote):

you mean the simplified one?

OpenStudy (lgbasallote):

4.1

OpenStudy (anonymous):

@rajathsbhat , is my formula incorrect?

OpenStudy (anonymous):

what's 10(2.1)^2 -10(2)^2 supposed to mean?

OpenStudy (lgbasallote):

uhmm substituting t = 2.1 and t = 2

OpenStudy (anonymous):

yes @Limitless .Absolutely

mathslover (mathslover):

that is the final answer

OpenStudy (anonymous):

Use my formula, Ig. Let \(t_1=2.1\) and \(t_2=2\).

OpenStudy (lgbasallote):

i wrote the follow-up problem lol..find average velocity for time interval from 2 to 2.10 s

OpenStudy (anonymous):

..I know

OpenStudy (anonymous):

It still applies.

OpenStudy (lgbasallote):

mehh i dont like formulas @_@

OpenStudy (lgbasallote):

heh i have blue eyes

mathslover (mathslover):

@lgbasallote see the images that i had uploaded ... that are the final answers..

OpenStudy (anonymous):

That's what math is about, lololol.

OpenStudy (lgbasallote):

where did 40 and 60 and 2 come from?

OpenStudy (lgbasallote):

oh lol it's just a different way of doing the problem?

mathslover (mathslover):

yes...

mathslover (mathslover):

and i think that is the easier way to solve that problem

OpenStudy (anonymous):

No. @mathslover is calculating velocity, not average velocity.

OpenStudy (anonymous):

Unless I'm missing something.

mathslover (mathslover):

@Limitless i am not calculating the velocity u can see the whole process there is no mistake there in the process .....for calculation of average velocity ...

OpenStudy (lgbasallote):

hmm but i think it would be better to just use this method to be uniform and not get confused...like i said im reluctant to formulas so too much methods gonna clog me up

OpenStudy (anonymous):

You used the differential definition of velocity, @mathslover.

OpenStudy (lgbasallote):

anyway i got \[\frac{4.1}{0.1} = 41 m/s\] for the follow up problem

mathslover (mathslover):

http://www.ux1.eiu.edu/~cfadd/1350/Hmwk/Ch02/Ch2.html

mathslover (mathslover):

i think there is the answer of limitless reply

mathslover (mathslover):

http://3mech2011-12.mtlsd.wikispaces.net/file/view/AP+motion+in+1D+Solutions.pdf see the 3rd question

mathslover (mathslover):

1) -50 m/s 2) 4.1 m/s

OpenStudy (lgbasallote):

there has been a lot of replies but still none answered me :C can you set aside the arguments later lol

OpenStudy (anonymous):

I think whatever @Limitless has said is true. Especially that equation.

OpenStudy (lgbasallote):

can someone show me thesolution for the follow up problem using the 10t^2 method thingy??? just so i know if i did it right

OpenStudy (anonymous):

Thank you. I have been eating chips the entire time, so I wasn't able to rigorously check answers.

OpenStudy (lgbasallote):

this thread is long and it's getting laggy so let's be direct to the point lol :P

OpenStudy (anonymous):

\[x = 10t^2 \implies v_x = 20t\] \[v_{avg}=\frac{\int_2^3 20tdt}{\int_2^3 dt} \]idk if it's right but should be

OpenStudy (lgbasallote):

nooo not the calculus way T_T

OpenStudy (anonymous):

lgba, there are two methods. This one: \[v_{avg}=\frac{\frac{x(t_1)}{t_1}-\frac{x(t_2)}{t_2}}{2}\] Then: \[ v_{avg}=\frac{1}{b-a}\int_{a}^{b}v(t) dt \]

OpenStudy (anonymous):

why not?